# Towards Reliable Connectivity: Measurement-Driven Assessment of Starlink and OneWeb Non-Terrestrial and 5G Terrestrial Networks

### v1

@@ IV-A GPS Traces
Analyzing the latency values for Op. A and Op. B in Fig. 5(a), it can be seen that the OWD is stable with uniform values between 15 and 25 ms throughout the different locations. This fact is especially remarkable in the case of the downlink. For the uplink OWD, although the values remain stable, there are several instances where latency increases to values between 100 and 120 ms (see Op. A). In any case, the low latency values obtained for terrestrial networks are to be expected given the redundancy of base stations in this area, where there are multiple cells covering the user terminal that can function as backup even if the main cell is not operating correctly or is congested. In the case of the uplink, the link budget can play a fundamental role in the occasional increase in latency, since the transmission power in the user equipment is certainly more restricted than the transmission power in the base station. Regarding the satellite operators, there are multiple coverage interruptions along the route where both uplink and downlink OWD exceed several seconds. Given the characteristics of the route, the buildings on the different avenues and streets act as sources of LoS obstruction, causing these interruptions. After an in-depth analysis of the measurements, it was observed that the main coverage interruptions occur when the vehicle turns at an intersection. This suggests that the change in the vehicle’s orientation during the turn changes the spatial distribution of the satellites from the user equipment perspective, which eventually causes a momentary loss of visibility with some of the satellites, resulting in a loss of connectivity for several seconds. In the case of throughput analysis, similar behavior to that shown by latency is found. Cellular operators show stable values around the targeted 5/5 Mbps in both downlink and uplink due to the high density of available cells. In the case of satellite operators, constant connection drops are observed, with throughput instantly falling from 5 Mbps to 0 Mbps, suggesting a sudden blockage of the LoS. It is still significant that these drops tend to occur after turns at intersections.
IV-B One Way Delay Statistics
@@
TABLE II: Best connectivity ranking based on three one-way delay thresholds (100 ms, 300 ms, and 500 ms) for several technologies and scenarios. 1st (
@@
In the suburban scenario (see Fig. 6(b)), there is a generalized statistical improvement in terrestrial and non-terrestrial solutions. While the median values are comparable between the urban and suburban scenarios, there is an improvement in network reliability reflected in the improved CEDF tail distribution. Specifically, the downlink OWD is reduced from values of 352 ms (Op. A) and 323 ms (Op. B) in the urban environment to 99 ms (Op. A) and 204 ms (Op. B) in the suburban environment for 99.99% reliability. Regarding OneWeb and Starlink, the loss of coverage probability decreases from 19%/21% (urban DL/UL) and 12%/17% (urban DL/UL) to 2%/2% (suburban DL/UL) and 3%/3% (suburban DL/UL), respectively. These results indicate an improvement in radio conditions that enhances connectivity across all four interfaces. In the case of cellular networks, the improvement in LoS conditions and lower building density enable improved network KPIs despite the lower number of base stations per km2. In the case of satellite networks, the lower number of obstacles on the horizon also leads to a notable improvement in LoS conditions and, consequently, connectivity. It is noteworthy that when the horizon is mostly clear, as it is the case for the suburban environment compared to the urban scenario, the performance between the two satellite networks tends to become more homogeneous. Therefore, the redundancy provided by the larger number of satellites in Starlink is no longer as relevant.
Finally, the forest scenario is the most challenging among all tested connectivity solutions, showing the worst results among all the scenarios studied. For cellular operators, it is possible to establish downlink connectivity, although with significant degradation in the distribution tails, e.g., 414 ms one-way delay (Op. A) and loss of connectivity (Op. B) at the P99.99 percentile. This behavior worsens even further in the uplink, with latencies exceeding 500 ms in 4% (Op. A) and 26% (Op. B) of the samples, respectively. These results are consistent with the previous discussion related to the one-way delay asymmetry due to the limitation of the uplink link budget, whose influence is even more noticeable in scenarios with heavy signal attenuation and few base stations, such as in the forest. In the satellite case, this scenario is predominantly affected by the NLoS condition due to the presence of vegetation along the horizon (see Fig. 2(c)). Therefore, the probability of complete coverage loss is up to 56%/60% (DL/UL) in OneWeb and up to 49%/49% (DL/UL) in Starlink. In this case, Starlink’s larger constellation is once again relevant for improving visibility compared to OneWeb.
@@ IV-C Throughput Statistics
Finally, Table IV presents the ranking of the best connectivity solution for all three scenarios based on three possible thresholds: 0 Mbps (i.e., null connectivity), 2 Mbps, and 4 Mbps. As a general note, cellular solutions tend to meet all three thresholds with probabilities of around 99.9%, except in the forest scenario, where the limited infrastructure available degrades connectivity, especially in Op. B. Focusing on satellite operators, Starlink tends to be the best option in the urban scenario for downlink. In terms of uplink, Starlink is the better option (90.8% Starlink vs. 86.5% OneWeb) for establishing connectivity in the link, i.e., P(TP UL >> 0 Mbps). Nonetheless, when higher uplink throughput is required, e.g., 2 Mbps (71.2% Starlink vs. 82.2% OneWeb) or 4 Mbps (63.2% Starlink vs. 77.7% OneWeb), OneWeb offers better results. This could be due to the larger antenna size in OneWeb (see Section III), thus allowing for better radio propagation conditions, i.e., antenna gain and transmission power in the uplink. These results support the discussion in the previous Section IV.B, where it was indicated that Starlink is generally a better option in terms of connectivity capacity in blockage-prone environments due to its higher satellite density. However, under the assumption of LoS between ground equipment and satellites, the performance of both satellite solutions tends to become more homogeneous, although with a slight performance advantage for OneWeb due to its larger antenna size. This hypothesis is also supported by the results in suburban and forest environments. In particular, in suburban environments where excellent LoS conditions are expected, the performance of both solutions is very similar, with OneWeb having a slight advantage over Starlink. However, in forest environments, where the probability of NLoS is exceptionally high, Starlink tends to be a better solution.
IV-D Technology Comparison and Discussion
@@
TABLE IV: Best connectivity ranking based on three throughput thresholds (0 Mbps, 2 Mbps, and 4 Mbps) for several technologies and scenarios. 1st (
@@ V Multi-Connectivity Evaluation
Taking advantage of the synchronization when forwarding packets across the network in the setup proposed in Section III, it is possible to develop the concept of multi-connectivity. Specifically, multi-connectivity involves the packet duplication and packet forwarding through two or more independent network interfaces. This duplication of packets brings two direct advantages: (i) larger network reliability, since in the event of packet loss on one of the interfaces, the alternative interface may still be able to forward that packet, and (ii) better KPI performance, since in terms of latency it will be reduced to that of the interface with the lowest delay, and in terms of throughput, this will be lower bounded by that of the interface with the highest available throughput. From a practical point of view, duplication can be carried out at the network layer by encapsulating IP packets with a sequence number and unique identifiers associated with each of the available interfaces. Similarly, this process is reversible on reception by decapsulating the IP packet, making it a completely transparent process at the application level. Note that when the same IP packet is received by a different interface, it can be discarded if the sequence number has already been received through an alternative interface. Applications implementing the described process have been successfully developed and tested at Aalborg University (AAU). These open-source tools are publicly available for general use [39].
Taking into account the connectivity options available in the measurement campaign presented throughout the work, three possible types of multi-connectivity are proposed: (i) cellular-cellular (Cell./Cell.) multi-connectivity between Op. A and Op. B, (ii) satellite-satellite (Sat./Sat.) multi-connectivity between OneWeb and Starlink, and (iii) cellular-satellite (Cell./Sat.) multi-connectivity between Op. A and Starlink. The first case (Cell./Cell.) takes advantage of the already dense deployment of cellular network base stations to maximize the joint coverage area of the infrastructure of Op. A and Op. B. The second case (Sat./Sat.) relies exclusively on satellite technology in the event of total failure of the terrestrial infrastructure. The combination of both satellite networks aims to take advantage of the satellites in both constellations to minimize spots without ground coverage. The third case (Cell./Sat.) simultaneously combines TN-NTN technology, thereby leveraging the complete independence of both communication systems to enhance the resilience of the communication channel. Tables V and VI, as well as Figs. 9 and 10, show the percentile, CEDF, and EDF statistics for the one-way delay and throughput KPIs in the three scenarios, i.e., urban, suburban and forest, for the three multi-connectivity cases described.
@@ VI Satellite Static Performance
While the analysis has thus far focused on satellite connectivity under mobility, it is equally critical to evaluate performance under static conditions in obstacle-rich settings, such as urban environments. Satellite operators are aware of this issue and recommend positioning ground user equipment in areas with full visibility towards the sky. Notwithstanding, this may not always be possible for logistical reasons. Therefore, in addition to the mobility study, we propose studying the satellite connectivity of Starlink and OneWeb operators in three static locations in the urban area under analysis. Fig. 11 shows a $360\degree$ panoramic view for the three locations, as well as the elevation profile of the obstacles blocking the horizon of the user equipment, where $\theta=0\degree$ corresponds to the horizon and $\theta=90\degree$ to the zenith. Each area has been characterized based on two parameters: (i) $\theta_{\max}$, which is the maximum elevation angle blocked on the horizon by some element in the environment, e.g., a building or a tree, and (ii) $\theta_{\mu}$, which is the average blocked elevation angle along the $360\degree$ azimuth profile.
@@
Fig. 12 shows the one-way delay CEDF in DL/UL for each static scenario in both satellite technologies. In each scenario, a time window is considered during which the user equipment has visibility of at least 10 satellites from the OneWeb constellation and several tens from the Starlink constellation over time, thus acquiring 600k UDP packet one-way delay values to estimate the visibility and connectivity conditions of the satellite link for each OneWeb and Starlink technologies. For Starlink, it can be observed that regardless of the scenario studied, it is always possible to transmit UDP packets with latencies below 500 ms, indicating no network outages and stable coverage despite the presence of obstacles with maximum elevations of up to $67\degree$. However, the same effect is not observed with OneWeb connectivity, where latency spikes exceeding a 500 ms threshold are experienced with some regularity, suggesting a temporary loss of visibility with the satellites in the constellation. In particular, in the first scenario (see Fig. 11(a)) with $\theta_{\max}=49\degree$ and $\theta_{\mu}=29\degree$, the threshold is exceeded with a probability of 0.04%/0.2% (DL/UL). In the second scenario (see Fig. 11(b)), where the presence of obstacles increases with $\theta_{\max}=61\degree$ and $\theta_{\mu}=39\degree$, the probability of exceeding the latency threshold also increases to 8.4%/10.6% (DL/UL). Finally, in the third scenario (see Fig. 11(c)), where the antenna is placed particularly close to a building with $\theta_{\max}=67\degree$ and $\theta_{\mu}=39\degree$, the probability of exceeding the threshold shoots up to 39.8%/44.1% (DL/UL). These results suggest a direct relation between the probability of signal blocking, denoted by the high latencies found in communications, and the $\theta_{\max}$ present in our measurement scenario. While Starlink is able to overcome signal blocking through satellite redundancy on the horizon, OneWeb cannot do it due to a much smaller constellation.

### v2

@@ IV-B One Way Delay Statistics
In the suburban scenario (see Fig. 7(b)), there is a generalized statistical improvement in terrestrial and non-terrestrial solutions. While the median values are comparable between the urban and suburban scenarios, there is an improvement in network reliability reflected in the improved CEDF tail distribution. Specifically, the downlink OWD is reduced from values of 352 ms (Op. A) and 323 ms (Op. B) in the urban environment to 99 ms (Op. A) and 204 ms (Op. B) in the suburban environment for 99.99% reliability. Regarding OneWeb and Starlink, the loss of coverage probability decreases from 19%/21% (urban DL/UL) and 12%/17% (urban DL/UL) to 2%/2% (suburban DL/UL) and 3%/3% (suburban DL/UL), respectively. Taking a closer look at Fig. 7(b), the OWD DL CEDF for OneWeb in the suburban scenario exhibits a bimodal behavior characterized by two latency clusters at around 55-60 ms and 95-100 ms, respectively. This behavior is consistent with handover dynamics between different ground stations, as previously reported in [36]. Similar bimodal latency patterns have been associated with changes in gateway routing, where traffic is dynamically forwarded through geographically distributed ground stations. In our measurements, this behavior can be attributed to the connection of OneWeb satellites to different European ground stations (see Section II), which leads to different latency clusters. Note that these patterns are not as noticeable in urban and rural environments, as they are dominated by the NLoS condition and signal blockage, which intrinsically modifies the latency distributions, generating long tails without a clear bimodal behavior. Overall, the previous results indicate an improvement in radio conditions that enhances connectivity across all four interfaces compared to the urban scenario. In the case of cellular networks, the improvement in LoS conditions and lower building density enable improved network KPIs despite the lower number of base stations per km2. In the case of satellite networks, the lower number of obstacles on the horizon also leads to a notable improvement in LoS conditions and, consequently, connectivity. It is noteworthy that when the horizon is mostly clear, as it is the case for the suburban environment compared to the urban scenario, the performance between the two satellite networks tends to become more homogeneous. Therefore, the redundancy provided by the larger number of satellites in Starlink is no longer as relevant.
TABLE II: Best connectivity ranking based on three one-way delay thresholds (100 ms, 300 ms, and 500 ms) for several technologies and scenarios.
1st (
@@
Finally, the forest scenario is the most challenging among all tested connectivity solutions, showing the worst results among all the scenarios studied. For cellular operators, it is possible to establish downlink connectivity, although with significant degradation in the distribution tails, e.g., 414 ms one-way delay (Op. A) and loss of connectivity (Op. B) at the P99.99 percentile. This behavior worsens even further in the uplink, with latencies exceeding 500 ms in 4% (Op. A) and 26% (Op. B) of the samples, respectively. These results are consistent with the previous discussion related to the one-way delay asymmetry due to the limitation of the uplink link budget, whose influence is even more noticeable in scenarios with heavy signal attenuation and few base stations, such as in the forest. These constraints suggest that under conditions in which we can establish connectivity, the system typically operates close to or beyond its uplink capacity when attempting to sustain the target 5 Mbps UL throughput (see Section III). This may lead to more conservative modulation and coding schemes at the MAC layers, increased HARQ retransmissions, and scheduling and buffering delays, which translate into significantly increased uplink one-way delay and, therefore, OWD asymmetry. In the satellite case, this scenario is predominantly affected by the NLoS condition due to the presence of vegetation along the horizon (see Fig. 2(c)). Therefore, the probability of complete coverage loss is up to 56%/60% (DL/UL) in OneWeb and up to 49%/49% (DL/UL) in Starlink. In this case, Starlink’s larger constellation is once again relevant for improving visibility compared to OneWeb.
@@
Concerning the comparison between the hardware equipment used for each satellite constellation, i.e., Kymeta Hawk U8 (OneWeb) vs. Standard Actuated Gen 2 (Starlink), as indicated in Section III, the larger OneWeb antenna is focused on enterprise market share, while the compact Starlink antenna is focused on the mass market consumer. Due to the different sizes, and assuming the same efficiency and frequency band, the difference in antenna gain $\Delta G$ can be approximated from the factor between areas as $10log_{10}(A_{Starlink}/A_{OneWeb})=10log_{10}(0.155\,\,\textrm{m}^{2}/0.801\,\,\textrm{m}^{2})=-7.1\,\,\textrm{dB}$. To determine whether this difference is significant for the results, the time traces and the CEDFs from the measurement campaign have been examined. In scenarios prone to signal blocking, such as urban and forest areas, latency behavior tends to exhibit a binary pattern: either the network operates under regular conditions with occasional latency spikes due to, for instance, periodic reconfiguration of satellite network links [47], or prolonged coverage interruptions are found with OWDs exceeding 500 ms and sustained packet losses indicating continuous loss of connectivity. Under these conditions, where the signal is prone to NLoS propagation in the satellite frequency bands, the difference in antenna gain becomes irrelevant due to the high attenuation of tens of decibels that occurs when obstacles obstruct the link. In these scenarios, Starlink performance is superior, which can be attributed to the higher satellite density that mitigates the probability of signal blockage (see Table II). In scenarios with excellent visibility where NLoS conditions are not as relevant, such as suburban environments, the performance of both constellations converges in the high latency regime at very similar values. According to the ranking in Table II, OneWeb ensures an OWD of less than 500 ms with a probability 1.2%-1.3% higher than Starlink. In this scenario, where visibility for both constellations is favorable and similar performance is obtained, this slight difference in favor of OneWeb may be linked to the user terminal design difference.
@@ IV-C Throughput Statistics
TABLE IV: Best connectivity ranking based on three throughput thresholds (0 Mbps, 2 Mbps, and 4 Mbps) for several technologies and scenarios.
1st (
@@
Finally, Table IV presents the ranking of the best connectivity solution for all three scenarios based on three possible thresholds: 0 Mbps (i.e., null connectivity), 2 Mbps, and 4 Mbps. As a general note, cellular solutions tend to meet all three thresholds with probabilities of around 99.9%, except in the forest scenario, where the limited infrastructure available degrades connectivity, especially in Op. B. Focusing on satellite operators, Starlink tends to be the best option in the urban scenario for downlink. In terms of uplink, Starlink is the better option (90.8% Starlink vs. 86.5% OneWeb) for establishing connectivity in the link, i.e., P(TP UL >> 0 Mbps). Nonetheless, when higher uplink throughput is required, e.g., 2 Mbps (71.2% Starlink vs. 82.2% OneWeb) or 4 Mbps (63.2% Starlink vs. 77.7% OneWeb), OneWeb offers better results. This may result from the larger antenna size in OneWeb (see Section III and IV.B), thus allowing for more beneficial radio propagation conditions, i.e., antenna gain and transmission power in the uplink. These results support the discussion in the previous Section IV.B, where it was indicated that Starlink is generally a better option in terms of connectivity capacity in blockage-prone environments due to its higher satellite density. However, under the assumption of LoS between ground equipment and satellites, the performance of both satellite solutions tends to become more homogeneous, although with a slight performance advantage for OneWeb due to its larger antenna size. This hypothesis is also supported by the results in suburban and forest environments. In particular, in suburban environments where excellent LoS conditions are expected, the performance of both solutions is very similar, with OneWeb having a slight advantage over Starlink. However, in forest environments, where the probability of NLoS is exceptionally high, Starlink tends to be a better solution.
@@ V Multi-Connectivity Evaluation
Taking advantage of the synchronization when forwarding packets across the network in the setup proposed in Section III, it is possible to develop the concept of multi-connectivity. Specifically, multi-connectivity involves the packet duplication and packet forwarding through two or more independent network interfaces. This duplication of packets brings two direct advantages: (i) larger network reliability, since in the event of packet loss on one of the interfaces, the alternative interface may still be able to forward that packet, and (ii) better KPI performance, since in terms of latency it will be reduced to that of the interface with the lowest delay, and in terms of throughput, this will be lower bounded by that of the interface with the highest available throughput. In exchange, the main cost of this approach lies in the redundancy of transmitted packets, which increases linearly with the number of interfaces used. Therefore, when two interfaces are considered, the traffic volume doubles. From a hardware perspective, the additional equipment must also be taken into account, as multi-connectivity implies multiple modems and/or user terminals, which leads to higher energy consumption and economic cost of the hardware and subscriptions to terrestrial and non-terrestrial networks. From a practical point of view, duplication can be carried out at the network layer by encapsulating IP packets with a sequence number and unique identifiers associated with each of the available interfaces. Similarly, this process is reversible on reception by decapsulating the IP packet, making it a completely transparent process at the application level. Note that when the same IP packet is received by a different interface, it can be discarded if the sequence number has already been received through an alternative interface. Applications implementing the described process have been successfully developed and tested at Aalborg University (AAU). These open-source tools are publicly available for general use [48]. In practice, the IP packet encapsulation involves an overhead of approximately 2.7% for 1500 Bytes IP packets. As indicated in Section III, multi-connectivity performance is evaluated using synchronized time traces acquired from all available interfaces during the measurement campaign. This approach enables a fair comparison between single-connectivity and multi-connectivity cases under identical radio and physical conditions, which is not possible with multiple drive tests using live multi-connectivity implementations, where only one solution can be assessed at a time. While this methodology does not aim to emulate the full control plane behavior of a specific live multi-connectivity protocol (e.g., packet reordering, buffering, or timeout mechanisms), it allows us to evaluate the performance gains achieved by multi-connectivity due to interface diversity and packet duplication. Experimentation in controlled and real environments has shown that the latency and throughput distributions obtained through post-processing synchronization closely match those observed in live multi-connectivity implementations, with performance overhead due to additional processing remaining negligible compared to end-to-end network delays [49].
@@
Taking into account the connectivity options available in the measurement campaign presented throughout the work, three possible types of multi-connectivity are proposed: (i) cellular-cellular (Cell./Cell.) multi-connectivity between Op. A and Op. B, (ii) satellite-satellite (Sat./Sat.) multi-connectivity between OneWeb and Starlink, and (iii) cellular-satellite (Cell./Sat.) multi-connectivity between Op. A and Starlink. The first case (Cell./Cell.) takes advantage of the already dense deployment of cellular network base stations to maximize the joint coverage area of the infrastructure of Op. A and Op. B. The second case (Sat./Sat.) relies exclusively on satellite technology in the event of total failure of the terrestrial infrastructure. The combination of both satellite networks aims to take advantage of the satellites in both constellations to minimize spots without ground coverage. The third case (Cell./Sat.) simultaneously combines TN-NTN technology, thereby leveraging the complete independence of both communication systems to enhance the resilience of the communication channel. Tables V and VI, as well as Figs. 10 and 11, show the percentile, CEDF, and EDF statistics for the one-way delay and throughput KPIs in the three scenarios, i.e., urban, suburban and forest, for the three multi-connectivity cases described.
@@ VI Satellite Static Performance
While the analysis has thus far focused on satellite connectivity under mobility, it is equally critical to evaluate performance under static conditions in obstacle-rich settings, such as urban environments. Satellite operators are aware of this issue and recommend positioning ground user equipment in areas with full visibility towards the sky. Notwithstanding, this may not always be possible for logistical reasons. Therefore, in addition to the mobility study, we propose studying the satellite connectivity of Starlink and OneWeb operators in three static locations in the urban area under analysis. Fig. 12 shows a $360\degree$ panoramic view for the three locations, as well as the elevation profile of the obstacles blocking the horizon of the user equipment, where $\theta=0\degree$ corresponds to the horizon and $\theta=90\degree$ to the zenith. The elevation profile was obtained by computing the elevation angle of the boundary curve separating street-level obstacles (NLoS) from the sky (LoS). Each area has been characterized based on two parameters: (i) $\theta_{\max}$, which is the maximum elevation angle blocked on the horizon by some element in the environment, e.g., a building or a tree. Therefore, $\theta_{max}$ is the maximum elevation angle attained over the full azimuthal range $\phi\in[0\degree,360\degree)$. (ii) $\theta_{\mu}$, which is the average blocked elevation angle along the $360\degree$ azimuth profile. It can be formally defined as $\theta_{\mu}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i}$, where $N$ denotes the number of points used to discretize the elevation profile along the azimuthal range.
@@
Fig. 13 shows the one-way delay CEDF in DL/UL for each static scenario in both satellite technologies. In each scenario, a time window is considered during which the user equipment has visibility of at least 10 satellites from the OneWeb constellation and several tens from the Starlink constellation over time, thus acquiring 600k UDP packet one-way delay values to estimate the visibility and connectivity conditions of the satellite link for each OneWeb and Starlink technologies. For Starlink, it can be observed that regardless of the scenario studied, it is always possible to transmit UDP packets with latencies below 500 ms, indicating no network outages and stable coverage despite the presence of obstacles with maximum elevations of up to $67\degree$. Note that the divergence observed at the tail of the CEDF in the uplink is attributable to the global network reconfiguration that Starlink performs every 15 seconds, leading to latency spikes when handovers occur between the ground station and satellite, and between the satellite and user terminal [47]. These spikes every 15 seconds have been confirmed by the time traces, thus ruling out the lack of visibility due to the presence of obstacles. Therefore, the behavior of the network in the three scenarios is stable according to Starlink’s procedures, regardless of the environment. However, the same effect is not observed with OneWeb connectivity, where latency spikes exceeding a 500 ms threshold are experienced with some regularity, suggesting a temporary loss of visibility with the satellites in the constellation. In particular, in the first scenario (see Fig. 12(a)) with $\theta_{\max}=49\degree$ and $\theta_{\mu}=29\degree$, the threshold is exceeded with a probability of 0.04%/0.2% (DL/UL). In the second scenario (see Fig. 12(b)), where the presence of obstacles increases with $\theta_{\max}=61\degree$ and $\theta_{\mu}=39\degree$, the probability of exceeding the latency threshold also increases to 8.4%/10.6% (DL/UL). Finally, in the third scenario (see Fig. 12(c)), where the antenna is placed particularly close to a building with $\theta_{\max}=67\degree$ and $\theta_{\mu}=39\degree$, the probability of exceeding the threshold shoots up to 39.8%/44.1% (DL/UL). These results suggest a direct relation between the probability of signal blocking, denoted by the high latencies found in communications, and the $\theta_{\max}$ present in our measurement scenario. While Starlink is able to overcome signal blocking through satellite redundancy on the horizon, OneWeb cannot do it due to a much smaller constellation.